how to convert 300 mm to m
Answer by sdmmadam@yahoo.com(530)

how to convert 300 mm to m
1000 mm = 1 m
Therefore 1 mm = (1/1000) of a metre
Therefore 300 mm = (1/1000) X (300)
= (3/10)
= 0.3 metre
Answer: 0.3 metre
Verification: 1 metre = 1000 mm
Therefore 0.3 metres = (1000)X(0.3) = 300.0 = 300 mm